OpenStudy (gabylovesyou):

Simplify 6(2r − 5) + 2r using the Distributive property

OpenStudy (anonymous):

Amistre just wait.

OpenStudy (amistre64):

the distributive property tells us we can multiply all the inner numbers in the paranthesis by the multiplier on the outside

OpenStudy (amistre64):

just wait for what?

OpenStudy (anonymous):

I think Gabby will give her answer to verify..

OpenStudy (gabylovesyou):

yes i am but give me a sec

OpenStudy (anonymous):

Okay take your time..

OpenStudy (amistre64):

my advice is still good tho :)

OpenStudy (gabylovesyou):

14r - 3????

OpenStudy (gabylovesyou):

How do i do it that way @amistre64 ??

OpenStudy (anonymous):

I though you will simplify it further.. but you did not.. Thanks.. gabby posts the question and then the answer solved by her and then she asks to verify..

OpenStudy (amistre64):

the 14r is good

OpenStudy (amistre64):

i think you typoed theeend; did you mean -3? or something else

OpenStudy (gabylovesyou):

These r my choices : 4r - 30 4r + 30 14r - 30 12r - 30

OpenStudy (amistre64):

then you most likely meant the -30

OpenStudy (anonymous):

what do you get 6*(-5)??

OpenStudy (gabylovesyou):

-30

OpenStudy (anonymous):

Yes so what is the answer??

OpenStudy (gabylovesyou):

14r - 30?! i think lol

OpenStudy (anonymous):

Yes you are right gabby..

OpenStudy (gabylovesyou):

yaay!! ok thanks :) (im not good at distributive property) :S

OpenStudy (anonymous):

See what distributive property says: \[\huge a(b+c) = a \times b + a \times c\]

OpenStudy (gabylovesyou):

but how come i got 12r - 30??

OpenStudy (gabylovesyou):

ohhhh i get it now

OpenStudy (gabylovesyou):

thanks :)

OpenStudy (anonymous):

yes you got right.. But there is 2r in the right also.. So. it became 12r + 2r = 14r. Welcome dear..

OpenStudy (gabylovesyou):

yes thats what i forgot :) What if it was 1/2(g - 16) would the answer be 1/2g - 8??

OpenStudy (gabylovesyou):

@waterineyes :)

OpenStudy (anonymous):

Yes you are right..

OpenStudy (gabylovesyou):

ok thanks :)

OpenStudy (anonymous):

Simple, no??

OpenStudy (gabylovesyou):

yea! Sometimes... lol i have 5 questions left in this hw assignment :P Im gonna try and then check it with u :)

OpenStudy (anonymous):

Sure..